perimeter of rectangle is 74 inches. What are dimensions if the length is 7 inches greater than twice the width.
Answer by nerdybill(7384)

perimeter of rectangle is 74 inches. What are dimensions if the length is 7 inches greater than twice the width.
.
Let w = width
then
2w+7 = length
.
2(w + 2w+7) = 74
2(3w+7) = 74
3w+7 = 37
3w = 30
w = 10 inches
.
Length:
2w+7 = 2(10)+7 = 20+7 = 27 inches
.
answer: 10 by 27 inches